In June ,Katherine bought 21 cans of food and 3 bags of liter for a total of $42 .in July she spent $60 on 32 cans of food and 4 bags of liters .What is the cost of one bag of liter?
Answer by jorel1380(3719)

21n+3p=42
32n+4p=60
84n+12p=168
96n+12p=180
12n=12
n=1
p=7
One bag of litter costs $7..